1.     Given TC = 100 + 60Q – 12Q2 + Q3, find:                                                

a.      The equations of the TVC, AVC, and MC functions.

b.     The level of output at which AVC and MC are minimum, and prove that the AVC and MC curves are U-shaped.

c.      Find the AVC and MC for the level of output at which the AVC curve is minimum.



1
Expert's answer
2021-01-19T07:45:59-0500

From the Total cost TC ;If Q=0,TC=100 and TFC=100

Therefore TVC=60Q-12Q2+Q3


AVC="\frac{TVC} {Q}" ="\frac{(60Q−12Q2+Q3)} {Q}"


AVC=60-12Q+Q2


MC="\frac{d(TC)} {dq}" ="\frac{d(60Q-12Q^2+Q^3)} {dq}"


MC=60-24Q+3Q2


b) For Minimum Output AVC ;"\frac{d(AVC)} {dq} =0"


"\frac{d(60-12Q+Q^2) }{dq} =0"


"\frac{d(AVC)} {dQ}" =-12 +2Q=0


Q=6

Hence AVC is minimum when output, Q=6


For Minimum output MC; 60-24Q+3Q2


"\frac{d(MC)} {dQ}" =-24+6Q=0


Q=4

Hence MC is minimum when output, Q=4


To prove that AVC and MC curves are U shaped ;


"\frac{d(MC)} {dQ}" =-24+6Q


"\frac{d2(MC)} {d2(Q)}" =6>0, hence the curves are U shaped.


c) AVC=60 - 12Q+Q2


60-12(6)+62

60-72+36

=24

The output level Q=6 at which AVC function is minimum, AVC and MC is 24.